OpenStudy (aleah54):

FAN AND MEDAL The first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. 1.Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer.

OpenStudy (aleah54):

@aaronq @AloneS @Atsie @Azionne @ace-n-it @Angiedoesschool2275 @batmano3 @bigpapa1 @briannaontimeforreal @bannana43 @Bobo-i-bo @baewolfstar @Conqueror @Cuanchi @countsrnc @chris215 @caseysorensen @CoinBank77 @countrygirl4life

OpenStudy (angiedoesschool2275):

Please do not tag multiple people, I am sorry I can not assist you but I am sure someone you have tagged can. I am in a rush to leave and I want you to be provided with the best help possible. My apologies! D:

OpenStudy (aleah54):

sorry and thank you for trying

OpenStudy (angiedoesschool2275):

My pleasure! This question seems like it may take a long time to assist you with since I prefer having no time limit at all so it allows me to do my best to help you. :3 Well, I have to go now but hopefully I can help you later on. Have a wonderful evening! ^-^

OpenStudy (aleah54):

Thank you and you too

OpenStudy (aleah54):

@Aveline @likeabossssssss

likeabossssssss (likeabossssssss):

Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. g(x) = x3 − x2 − 4x + 4 g(x) = x3 + 2x2 − 9x − 18 g(x) = x3 − 3x2 − 4x + 12 g(x) = x3 + 2x2 − 25x − 50 g(x) = 2x3 + 14x2 − 2x − 14

likeabossssssss (likeabossssssss):

medal?